A proton accelerates from rest in a uniform electric field of 600 N/C. At one later moment, its speed is 1.30 Mm/s (nonrelativistic because v is much less than the speed of light) (a) Find the acceleration of the proton 5.74e10m s2 (b) Over what time interval does the proton reach this speed? Your response is within 10% of the correct value. This may be due to roundoff error, or you could have a mistake in your calculation, carry out a目intermediate results to at least four-digit accuracy to minimize roundoff errors c How far does it move in this time interval? 2.38 Your response differs from the correct answer by more than 10%. Double check your calcula ons m (d) What is its kinetic energy at the end of this interval? 1.41e-33 Our response differs significantly from the correct answer Rework your solution from the beginning and check each step carefuly.
Answer #1

b) using equation of the motion

V=u+at ...1

here u=0 (from rest), V=1.3×10^6 m/s,a=5.74×10^10 m/s^2

From equation 1

t=V-u/a

t=1.3×10^6/5.74×10^10

t=0.2265×10^(-4)

t=2.265×10^(-5) s. Answer

c) using equation of the motion

S=ut+at^2/2

here u=0 , a=5.74×10^10 m/s^2, t=2.265×10^(-5)s

S=0+5.74×10^10×(2.265×10^(-5))^2/2

S=14.72 m. Answer

d)

kinetic energy =MV^2/2

here M=1.673×10^(-27) , V=1.3×10^6 m/s

K= 1.673×10^(-27)×(1.3×10^6)^2 /2

K=1.414 ×10^(-15) J. Answer